OpenStudy (anonymous):

solve A=1/2h(b+B)for h

OpenStudy (anonymous):

A = 1/2h(b+B) for h ; multiply both sides by 2 2A = h(b+B) ; divide both sides by quantity (b+B) 2A / (b+B) = h

OpenStudy (anonymous):

nj is right

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

the last line yes

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

What's confusing, I'll try my best to explain it! :)

OpenStudy (anonymous):

how the answer is written

OpenStudy (anonymous):

how do i graph that

OpenStudy (anonymous):

I'm not sure actually, I would think you'd have some values to input. You can't construct a graph of variables. Did anything come with the problem?

OpenStudy (anonymous):

graphing that would be like 4d

OpenStudy (anonymous):

i meaN I DONT HAVE TO GRAPH

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

haha. If it's any concellation, your original equation is the area of a trapezoid xD